OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively.

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_2=-2}\) \(\large\color{blue}{ a_5=16}\) At first to find the ratio, \(\large\color{blue}{ a_n=a_{n-v} \times r^{v}}\) in general, and in your case, \(\large\color{blue}{ a_5=a_{2} \times r^{5-3}}\)

OpenStudy (solomonzelman):

After that find \(\large\color{blue}{ a_1}\) then you can fill in the \(\large\color{black}{ a_n=a_1 \times r^{n-1}}\)

OpenStudy (solomonzelman):

let me know if you need help

OpenStudy (anonymous):

how do i find a1?

OpenStudy (anonymous):

is the answer an = 1 • (-2)^(n - 1)

OpenStudy (anonymous):

-2 would be the second term and 16 would be the 5th term

OpenStudy (solomonzelman):

Have you found the common ratio ?

OpenStudy (anonymous):

would it be -8?

OpenStudy (anonymous):

do i have to divide by 2?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ a_5 =a_2 \times r^{5-2} }\) \(\Large\color{blue}{ 16 =(-2)\times r^{5-2} }\) \(\Large\color{blue}{ -8 = r^{5-2} }\) \(\Large\color{blue}{ -8 = r^{3} }\) \(\Large\color{blue}{ (-2)^3 = r^{3} }\) \(\Large\color{blue}{ r=? }\)

OpenStudy (anonymous):

-2

OpenStudy (solomonzelman):

Yes, there we go !

OpenStudy (solomonzelman):

Can you find a1 ?

OpenStudy (anonymous):

would it be 1?

OpenStudy (solomonzelman):

Yes !

OpenStudy (solomonzelman):

So you know that a1=1 r=-2 fill in the equation for any nth term below. \(\Large\color{blue}{ a_n=a_1\times r^{n-1} }\)

OpenStudy (anonymous):

an=1*(-2)^(n-1)

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

Do you want to type × like me ?

OpenStudy (solomonzelman):

multiplication symbol shortcut , no ?

OpenStudy (anonymous):

how?

OpenStudy (solomonzelman):

1) Click and hold ALT 2) click the number code (using the numbers that are on the right of the keyboard, and `NOT` the ones below `F1`, `F2`, `F3`, etc., ) 3) release the ALT number code result `0 2 1 5 ` × `2 4 6 ` ÷ `7 5 4 ` ≥ `7 5 5` ≤ `2 4 1 ` or `7 5 3` ± `2 4 7` ≈

OpenStudy (solomonzelman):

these is just a small part of it

OpenStudy (solomonzelman):

I'll stop for now

OpenStudy (anonymous):

0215

OpenStudy (solomonzelman):

No, CLICK AND HOLD ALT CLICK 0 2 1 5 on the number pad on the bottom right corner of your keyboard RELEASE ALT NOW

OpenStudy (anonymous):

oh ×

OpenStudy (solomonzelman):

YEs !

OpenStudy (anonymous):

Heh heh

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Anytime:)

OpenStudy (solomonzelman):

°¹²³ counting up to 3

OpenStudy (solomonzelman):

have fun !

OpenStudy (anonymous):

do you have time for more?

OpenStudy (solomonzelman):

more questions? I think so :)